A parallel plate air capacitor has a capacitance of 5 F. What is the electric field between two parallel metal plates separated by 2,5 cm with a potential difference of 3000V?

1
Expert's answer
2022-03-23T07:16:15-0400

The electric field between plates of capacitor

"E=\frac{V}{d}\\
=\rm\frac{3000\: V}{0.025\: m}=1.2\times 10^5\: V/m"